kekeman:

At which values in the interval [0, 2π) will the functions f (x) = cos 2x - 5 and g(x) = sin3x - 5 intersect? A. x = 0, π/10, π/2, 9π/10, 13π/10 B. x = 3π/10, π/2, 7π/10, 13π/10, 17π/10 C. x = π/10, π/2, 9π/10, 13π/10, 17π/10 D. x = 0, π/2, π/10, 13π/10, 17π/10

kekeman:

I thinking option B.)

Aiden3j:

Yep that should be correct

kekeman:

You sure?

kekeman:

@surjithayer

Aiden3j:

let me check

Aiden3j:

yep seems right u can still check with someone else first

kekeman:

Okay thanks

surjithayer:

\[\cos 2x-5=\sin 3x-5\] \[\cos 2x=\sin 3x\] \[\cos 2x=\cos (\frac{ \pi }{ 2 }-3x)\] \[\cos 2x-\cos (\frac{ \pi }{ 2 }-3x)=0\] \[\cos C-\cos D=-2\sin \frac{ C+D }{ 2 }\sin \frac{ C-D }{2 }\] \[-2\sin \frac{ 2x+\frac{ \pi }{ 2 } -3x}{ 2}\sin \frac{ 2x-\frac{ \pi }{ 2 } +3x}{ 2} =0\] either \[\sin (\frac{ \pi }{ 4 }-\frac{ x }{ 2 })=0=\sin (n \pi)\] \[\frac{ \pi }{ 4 }-\frac{ x }{ 2 }=n \pi \] \[\pi-2x=4n \pi \] \[2x=\pi-4n \pi \] n=0 \[x=\frac{ \pi }{ 2 }\] n=-1 \[x=\frac{ 5\pi }{ 2 },rejected\] \[x \in[0,2\pi]\] or \[\sin (\frac{ 5x }{ 2 }-\frac{ \pi }{ 4 })=0=\sin n \pi \] \[\frac{ 5x }{ 2 }-\frac{ \pi }{ 4 }=n \pi,n\in I\] \[10x-\pi=4n \pi\]\[10x=4n \pi+\pi \] n=0 \[x=\frac{ \pi }{ 10 }\] n=1 \[x=\frac{ 5\pi }{ 10 }=\frac{ \pi }{ 2 }\] n=2 \[10x=9\pi \]\[x=\frac{ 9\pi }{ 10}\] n=3 \[10x=13\pi \]\[x=\frac{ 13 \pi }{ 10 }\] n=4 \[10x=17\pi \]\[x=\frac{ 17\pi }{ 10}\] n=5 \[10x=20\pi+\pi=21\pi\]\[x=\frac{ 21\pi }{ 10 }\]\[Rejected\]

surjithayer:

so C

kekeman:

Ohhh I see wow thank you sm!